Solve the problem


A Ferris wheel with radius 15.0 m is turning about a horizontal axis through its center; the linear speed of the passenger rim is constant and equal to 9 m/s.





What are the magnitude and direction of the acceleration of a passenger as she passes through the lowest point in her circular motion?





How much time does it take the Ferris wheel to make one revolution?

1
Expert's answer
2021-10-31T18:09:43-0400
"a=\frac{9^2}{15}=5.4\frac{m}{s^2}"

Upwards.


"t=\frac{2\pi (15)}{9}= 10.5\ s"